A manufacturer of tobacco products plans to market a new brand of cigarette. The regulatory
commission needs to know the mean tar content for the new brand. Analysis of a random sample
of 25 of the new brand of cigarettes gives a mean tar content of 10.98 milligrams with a standard
deviation of 0.604 milligram. Construct a 95% confidence interval for the mean tar content of
the new brand.
1
Expert's answer
2021-10-04T13:48:58-0400

Population standard deviation is unknown , we will use t test

Degrees of freedom "=n-1=25-1=24"

Critical value=

"t_{\u03b1\/2, df} = t_{0.05\/2.24} = t_{0.025, 24} = \u00b12.0639"

95% confidence interval is given by

"\bar{x} ± t_{α/2, df} \times \frac{s}{\sqrt{n}} \\

= (10.98 -2.0639 \times \frac{0.604}{\sqrt{25}} < \mu < 10.98 + 2.0639 \times \frac{0.604}{\sqrt{25}}) \\

= (10.98 -0.24932 < \mu < 10.98 + 0.24932) \\

= (10.73068 < \mu < 11.22932)"